### v1

# Partial Functional Dynamic Backdoor Diffusion-based Causal Model

Abstract
We introduce a Partial Functional Dynamic Backdoor Diffusion-based Causal Model (PFD-BDCM), specifically designed for causal inference in the presence of unmeasured confounders with spatial heterogeneity and temporal dependency. The proposed PFD-BDCM framework addresses the restrictions of the existing approaches by uniquely integrating models for complex spatio-temporal dynamics with the analysis of multi-resolution variables. Specifically, the framework systematically mitigates confounding bias by integrating valid backdoor adjustment sets into a diffusion-based sampling mechanism. Moreover, it accounts for the intricate dynamics of unmeasured confounders through the deployment of region-specific structural equations and conditional autoregressive processes, and accommodates variables observed at heterogeneous resolutions via basis expansions for functional data. Our theoretical analysis establishes error bounds for counterfactual estimates of PFD-BDCM, formally linking reconstruction accuracy to counterfactual fidelity under monotonicity assumptions of structural equation and invertibility assumptions of encoding function. Empirical evaluations on synthetic datasets and real-world air pollution data demonstrate PFD-BDCM’s superiority over existing methods.
Keywords: Unmeasured Confounders; Spatiotemporal Dynamic Causal Graphs; Diffusion-based Causal Models; Backdoor Criterion
1 Introduction
Causal inference fundamentally addresses interventional (“What if?”) and counterfactual (“Why?”) questions that go beyond statistical correlations, proving valuable in high-stakes domains like healthcare for treatment effect estimation (Hill,, 2011), policy evaluation without randomized trials (LaLonde,, 1986). The field’s core challenge stems from the fundamental problem of causal inference: the impossibility of simultaneously observing both an outcome under a treatment and the potential outcomes under the alternative treatment (or control) for the same unit (Imbens and Rubin,, 2015), necessitating methods to overcome confounding bias in observational data. Traditional approaches including potential outcomes frameworks (Imbens and Rubin,, 2015), propensity scoring (Rosenbaum and Rubin,, 1983), instrumental variables (Angrist et al.,, 1996), and structural causal models (Pearl,, 2009) exhibit significant limitations when handling modern complex datasets—they struggle with high-dimensional confounders, ethical constraints of randomized trials, scarcity of valid instruments, and requirement of known causal graphs. These limitations become particularly acute when confounders involve high-dimensional data like medical images or partially unobserved variables (Shalit et al.,, 2017).
In the Structural Causal Model (SCM) framework, causal queries can be answered by learning a proxy for the unobserved exogenous noise and the structural equations (Pearl,, 2009). This suggests that (conditional) generative models that encode to a latent space could be a option for modeling SCMs, as the latent space serves as proxies for exogenous noises. In the SCM, the encoding process extracts the latent space from an observation, and the decoding process generates the sample from the latent space, approximating the structural equations. Chao et al., (2023) generalized their approach to arbitrary causal graphs, and proposed the Diffusion-based Causal Model (DCM) that allowed one to sample from the target distribution of interest, by which it can estimate the average treatment effect (ATE), outperforming some of the state-of-the-art algorithms (Sanchez-Martin et al.,, 2021; Khemakhem et al.,, 2021). DCM assumes causal sufficiency (i.e., absence of hidden confounders) and focused on approximating the structural equations given observational data and the underlying causal directed acyclic graph (DAG), and providing a mechanism for answering the causal observational, interventional and counterfactual queries (Pearl,, 2009). The DCM assumes causal sufficiency, which needs observing all confounding variables, which usually fails in applications. To overcome the limitation, Shimizu, (2023) proposed the BDCM algorithm for estimating ATE in the presence of unmeasured confounders by including nodes that satisfy the backdoor criterion (Pearl et al.,, 2016) in both training and sampling phases.
Both DCM and BDCM are for static causal relationships, whereas real-world causal relationships among variables may be affected by the spatio-temporal unmeasured confounding factors. Therefore, we propose a Partial Functional Dynamic Backdoor Diffusion-based Causal Model (PFD-BDCM), which can uncover the causal relationships among variables across different resolutions in the presence of spatio-temporal unmeasured confounders.
Our Contributions.
We propose spatio-temporal diffusion models for modeling partial functional spatio-temporal dynamic causal relationships. Diffusion models (Sohl-Dickstein et al.,, 2015; Ho et al.,, 2020; Song et al.,, 2021) have gained prominence due to their high expressiveness and performance in generative tasks (Saharia et al.,, 2022; Ramesh et al.,, 2022; Kong et al.,, 2020). Our primary contribution is to show how to apply diffusion models to capture partially functional dynamic causal relationship with unmeasured confounders. The core idea is in modeling each node in dynamic causal graphs with a spatio-temporal diffusion model and cascading the generated samples in the topological order to answer causal queries. For each node, its corresponding spatio-temporal diffusion model encodes and decodes latent representations using variables of nodes. Within our framework, the encoding and decoding procedures extend the Denoising Diffusion Implicit Models (DDIMs) (Song et al.,, 2021) paradigm by integrating backdoor adjustment sets as supplementary covariates. We term the resulting framework as the Partial Functional Dynamic Backdoor Diffusion-based Causal Model (PFD-BDCM). Theoretical analysis demonstrate that PFD-BDCM preserves essential structural properties of Structural Causal Models (SCMs). Key contributions of our study include:
  1. 1.
    The proposed PFD-BDCM provides a unified framework for approximating both interventions (do-operator) and counterfactuals (abduction-action-prediction steps). It has a training procedure requiring only the dynamic causal graph and observational data, and the trained model enables: i) sampling from observational/interventional distributions; ii) precise counterfactual query resolution.
  2. 2.
    Our theoretical analysis proves that the counterfactual estimates given by PFD-BDCM admit quantifiable error bounds under reasonable assumptions. i) It provides the first formal explanation for the performance gains of encoder-decoder architectures (e.g., diffusion models) in counterfactual querying through error bounds; ii) It extends to the more challenging multivariate case under an additional assumption and to diverse encoder-decoder models.
  3. 3.
    We evaluate the performance of PFD-BDCM in empirical settings involving spatio-temporal dynamic structure equations and three forms of causal queries.
2 Preliminaries
Notations.
To distinguish between the nodes in the causal graph and diffusion random variables, we use subscripts to denote graph nodes. Let $n]:=\{1,\cdots,n\}$, $\mathrm{dim(x)}$ represents the dimension of $x$.
Structural Causal Models.
Consider a directed acyclic graph (DAG) $\mathcal{G}$ with nodes $[K]$ in a topologically sorted order  [^1] A topological order is a linear arrangement of variables where a variable appears after all its direct causes (parents) (Pearl,, 2009)., where a node $k$ is built with an endogenous random variable $X_{k}$ defined on a space $\mathcal{X}_{k}\subset\mathbb{R}^{d_{k}}$, which has a random exogenous input $U_{k}$. Let $\mathrm{PA}_{k}$ be the set of parent nodes of node $k$ in $\mathcal{G}$ and let $X_{\mathrm{PA}_{k}}:=\left\{X_{l}\right\}_{l\in\mathrm{PA}_{k}}$ be a set of parent random variables defined on $\mathrm{PA}_{k}$.
A structural causal model (SCM) $\mathcal{M}$ characterizes the relationship between the endogenous variable of a node with the endogenous variables of its parents and its own exogenous variable. Formally, we define $\mathcal{M}:=(\mathbf{F}(\mathbf{X},\mathbf{U}),p_{U})$, where $\mathbf{F}(\mathbf{X},\mathbf{U})$ specifies how entire endogenous variables $\mathbf{X}:=\{X_{1},\cdots,X_{K}\}$ are generated from the set of exogenous random variables $\mathbf{U}:=\{U_{1},\cdots,U_{K}\}$ with a prior distribution $p_{U}$. The structural mechanism is governed by $\mathbf{F}(\mathbf{X},\mathbf{U}):=(f(X_{\mathrm{PA}_{1}},U_{1}),\cdots,f(X_{\mathrm{PA}_{K}},U_{K}))$, where each $X_{k}:=f(X_{\mathrm{PA}_{k}},U_{k})$ for $k\in[K]$ (Pearl,, 2009).
Conditional Diffusion Model.
Recent advances in deep learning-based causal inference (Chao et al.,, 2023; Shimizu,, 2023) have demonstrated the effectiveness of diffusion models for answering causal queries. Our approach leverages the generative power of diffusion models to learn the complex functional relationships inherent in a Structural Causal Model. The core idea is to represent the structural equation for each endogenous variable $X_{k}$ with a dedicated conditional diffusion model. This model learns the distribution $p(X_{k}|X_{\mathrm{PA}_{k}})$ .
Diffusion models (Sohl-Dickstein et al.,, 2015; Ho et al.,, 2020) approximate a target data distribution $q(x^{0})$ via a two-stage process. First, a fixed forward process gradually injects Gaussian noise into the data $x^{0}$ over $T$ steps. The distribution of the noisy data $x^{t}$ at step $t$ is a Gaussian: $q(x^{t}|x^{0})=\varphi(x^{t};\sqrt{\alpha_{t}}x^{0},(1-\alpha_{t})I)$, where ${\varphi}(x;\mu,\Sigma)$ denote the Gaussian density with mean $\mu$ and covariance $\Sigma$. Here, $\alpha_{t}:=\prod_{s=1}^{t}(1-\beta_{s})$ where $\beta_{s}$ is a predefined noise schedule at each time step $s$. As $t\to T$, $x^{T}$ converges to a standard Gaussian distribution.
Second, a learnable reverse process, parameterized by $\theta$, is trained to denoise the data. In modern implementations, this is achieved by training a neural network $\epsilon_{\theta}$ to predict the added noise $\epsilon$, where $\theta$ represents learnable parameters of the neural network, conditioned on the noisy data $x^{t}$, the step $t$, and any contextual information $c$. The objective function is:
$\mathbb{E}_{t,x^{0},c,\epsilon}\left[\|\epsilon-\epsilon_{\theta}(\sqrt{\alpha_{t}}x^{0}+\sqrt{1-\alpha_{t}}\epsilon,c,t)\|^{2}\right].$ (1)
In DCM framework, the conditioning context $c$ for a variable $X_{k}$ is its set of parent variables $X_{\mathrm{PA}_{k}}$.
3 Partial Functional Dynamic Backdoor Diffusion-based Causal Model
In SCMs, the presence of unmeasured confounders exhibiting spatial heterogeneity and temporal dependencies renders existing DCMs and BDCMs inadequate for addressing genuine causal inquiries among variables. To surmount this fundamental limitation, we propose a Spatio-Temporal Dynamic Structural Causal Model (ST-DSCM) based on the Backdoor Criterion. We commence by establishing several essential definitions. Due to the complexity of the model, some symbols are mixed up in the article, but the overall readability remains unaffected.
In causal inference, a confounder denotes a variable that causally influences both a treatment variable $X$ and an outcome variable $Y$, thereby inducing a non-causal association between them. Formally, a confounder $Z$ must satisfy: i) $Z$ has a causal effect on $X$; ii) $Z$ has a causal effect on $Y$; iii) $Z$ is not affected by $X$ (i.e., not on the causal pathway between $X$ and $Y$) (Pearl,, 2009).
Observable Confounders refer to measurable variables that satisfy the above conditions and can be measured, which permit adjustment through statistical methods such as stratification, matching or regression (Pearl,, 2009).
Unobservable Confounders denote latent variables that fulfill confounding criteria but resist direct measurement, potentially biasing causal estimates when unaccounted for.
Unobserved explanatory (explained) nodes are unobserved confounder nodes that have no unobserved confounder nodes as its parent (descendant).
3.1 Spatio-Temporal Dynamic Structural Causal Model (ST-DSCM)
Figure 1: Spatio-Temporal Dynamic Structural Causal Model with twelve exogenous and endogenous nodes (where nodes $X_{5}$, $X_{6}$ and $X_{7}$ are unmeasured confounder variables with spatial heterogeneity and temporal dependencies)
Suppose there is a spatio-temporal dataset $\boldsymbol{X}:=\{X_{k}\}_{k\in[K]}$ containing $K$ variables across $n$ regions over $J$ time points. Using $i$ to index regions and $j$ for time points, $X_{k}=(X_{ijk})_{n\times J}$ represent the value of the $k$-th variable $X_{k}$ over the time points and regions. We consider a DAG to characterize the causal relationships among $\{X_{k}\}_{k\in[K]}$.
Consider a DAG $\mathcal{G}$ (e.g. Figure 1) with nodes $[K]$ and a topologically sorted order such that each node $k$ has the $X_{k}$ as the random variable. Let $\mathcal{B}_{k}$ be the set of backdoor nodes [^2] A set of node $\mathcal{B}$ satisfies backdoor criterion (Pearl et al.,, 2016) for tuple $(X,Y)$ in DAG $\mathcal{G}$ if no node in $\mathcal{B}$ is a descendant of $X$ and $\mathcal{B}$ blocks all paths between $X$ (cause) and $Y$ (outcome) which contains an arrow into $X$. of $k$, and $\boldsymbol{X}_{\mathcal{B}_{k}}:=\{X_{l}\}_{l\in\mathcal{B}_{k}}$ represent the variables on $\mathcal{B}_{k}$. Furthermore, let $\mathcal{C}_{1}$ and $\mathcal{C}_{2}\subseteq[K]$ be two distinct sets of nodes with unobserved confounders, where $\mathcal{C}_{1}$ designates a set of unobserved explanatory nodes and $\mathcal{C}_{2}$ denotes a set of unobserved explained nodes. For $h=1,2$, let $\boldsymbol{X}_{\mathcal{C}_{h}}:=\{\boldsymbol{X}_{{\mathcal{C}_{h}}ij}\}_{i\in[n],j\in[J]}:=\{X_{qij}\}_{i\in[n],j\in[J],q\in\mathcal{C}_{h}}$, $\boldsymbol{U}_{\mathcal{C}_{h}}:=\{\boldsymbol{U}_{{\mathcal{C}_{h}}ij}\}_{i\in[n],j\in[J]}:=\{U_{qij}\}_{i\in[n],j\in[J],q\in\mathcal{C}_{h}}$. For example in Figure  1, $\boldsymbol{X}_{\mathcal{C}_{1}}=(X_{5},X_{6})^{\top}$, $\boldsymbol{X}_{\mathcal{C}_{2}}=(X_{7})^{\top}$.
To incorporate spatio-temporal dynamic structures among unmeasured confounders, and to account for spatial heterogeneity, we posit a relationship such that
$\boldsymbol{X}_{{\mathcal{C}_{2}}ij}=\boldsymbol{\Gamma}_{i}\boldsymbol{X}_{{\mathcal{C}_{1}}ij}+\boldsymbol{U}_{{\mathcal{C}_{2}}ij},$ (2)
where $\boldsymbol{\Gamma}_{i}$ is a $\dim(\boldsymbol{X}_{{\mathcal{C}_{2}}ij})\times\dim(\boldsymbol{X}_{{\mathcal{C}_{1}}ij})$ structural coefficient matrix, which are permitted to vary across regions.
To capture the temporal dependence, we let $\boldsymbol{X}_{{\mathcal{C}_{1}}i}=(\boldsymbol{X}_{{\mathcal{C}_{1}}i1}^{\top},\cdots,\boldsymbol{X}_{{\mathcal{C}_{1}}iJ}^{\top})^{\top}$ and $\boldsymbol{U}_{{\mathcal{C}_{2}}i}=(\boldsymbol{U}_{{\mathcal{C}_{2}}i1}^{\top},\cdots,\boldsymbol{U}_{{\mathcal{C}_{2}}iJ}^{\top})^{\top}$, then, the covariance matrices $\boldsymbol{\Sigma}_{\boldsymbol{X}_{{\mathcal{C}_{1}}i}}$ and $\boldsymbol{\Sigma}_{\boldsymbol{U}_{{\mathcal{C}_{2}}i}}$ of $\boldsymbol{X}_{{\mathcal{C}_{1}}i}$ and $\boldsymbol{U}_{{\mathcal{C}_{2}}i}$ can be expressed as
$\boldsymbol{\Sigma}_{\boldsymbol{X}_{{\mathcal{C}_{1}}i}}=\boldsymbol{D}_{{\mathcal{C}_{1}}i}\otimes\boldsymbol{T}_{\boldsymbol{X}_{{\mathcal{C}_{1}}}},\ \boldsymbol{\Sigma}_{\boldsymbol{U}_{{\mathcal{C}_{2}}i}}=\boldsymbol{D}_{{\mathcal{C}_{2}}i}\otimes\boldsymbol{T}_{\boldsymbol{U}_{\mathcal{C}_{2}}},$ (3)
where $\boldsymbol{D}_{{\mathcal{C}_{h}}i}(h\in\{1,2\})$ are the $J\times J$ adjacent time covariance matrices, $\boldsymbol{T}_{\boldsymbol{X}_{{\mathcal{C}_{1}}}}$ and $\boldsymbol{T}_{\boldsymbol{U}_{{\mathcal{C}_{2}}}}$ represent between-variable covariances of $\boldsymbol{X}_{{\mathcal{C}_{1}}ij}$ and $\boldsymbol{U}_{{\mathcal{C}_{2}}ij}$, respectively.
By incooperating the spatio-temporal dynamic structure among unmeasured confounders into the traditional SCM framework, we develop a novel Spatio-Temporal Dynamic Structural Causal Model (ST-DSCM) to describe the relationship between an observed node $k$ and its causal backdoor nodes $\boldsymbol{X}_{\mathcal{B}}$ with unmeasured confounders $\{\boldsymbol{X}_{\mathcal{C}_{h}}\}_{h\in\{1,2\}}$ and $\boldsymbol{U}_{\mathcal{C}_{2}}$ exhibit spatio-temporal dynamic structural relationships as constructed in the above.
However, real-world applications often present significant challenges in elucidating causal relationships among variables observed under heterogeneity. To address these complexities, we extend the ST-DSCM by incorporating functional random variables $\boldsymbol{X}_{k}(t)$, leading to a Partial Functional Spatio-Temporal Dynamic Structural Causal Model (PFST-DSCM), as described in Figure 2.
3.2 PFST-DSCM: Partial Functional Spatio-Temporal Dynamic Structural Causal Model
We adopt a basis expansion framework to achieve dimensionality reduction in the functional space (Zoh et al.,, 2024) via a set of orthogonal basis functions $\{\mathbf{b}_{ij1},\cdots,\mathbf{b}_{ijK_{n}}\}\in\mathbb{R}^{T\times K_{n}}$. Let
$X_{ijm}=\int\mathbf{b}_{ijm}(t)X_{ij}(t)dt,$ (4)
for $i\in[n],j\in[J],m\in[K_{n}]$, which are used as nodes within the ST-DSCM (see Figure 2).
Figure 2: Partial Functional Spatio-Temporal Dynamic Structural Causal Model with 33 exogenous and endogenous nodes (where nodes $X_{28}$, $X_{29}$ and $X_{30}$ are unmeasured confounders with spatial heterogeneity and temporal dependencies, $Y_{1}(t),Y_{2}(t),Y_{3}(t)$ are functional nodes, and $X_{4},\cdots,X_{21}$ are the corresponding base expansion nodes)
We assume that the unobserved random variables are jointly independent (Markovian SCM), and the Spatio-Temporal Dynamic Directed Acyclic Graph (ST-DDAG) $\mathcal{G}$ is the graph induced by ST-DSCM $\mathcal{M}$. Every ST-DSCM $\mathcal{M}$ entails a unique joint observational distribution satisfying the causal Markov assumption: $p(\mathbf{X})=\prod_{k=1}^{K}p(X_{k}|X_{\mathcal{B}_{k}})$.
3.3 PFD-BDCM
We introduce the PFD-BDCM for modeling PFST-DSCM and addressing causal queries. The PFD-BDCM approach leverages on an encoder-decoder architecture to construct a PFST-DSCM, which facilitates elucidation of causal relationships among variables observed at varying resolutions in the presence of unmeasured confounders that exhibit spatial heterogeneity and temporal dependencies.
When unmeasured confounders exist with spatial heterogeneity and temporal dependencies, we model a data generating process for observed variables $\{\mathbf{X}_{k}:X_{ijk}=f_{ij}(\mathbf{X}_{\mathcal{B}_{k}},\mathbf{U}_{k})\}_{k\in[K]}$.
Our framework has an encoding function $g$ and a decoding function $h$. The encoder $g$ maps the observed data $(\mathbf{X}_{k},\mathbf{X}_{\mathcal{B}_{k}})$ to a latent variable $\mathbf{Z}_{k}$, defined as $\mathbf{Z}_{k}:=g(\mathbf{X}_{k},\mathbf{X}_{\mathcal{B}_{k}})$, aiming to encapsulate information pertaining to $\mathbf{U}_{k}$. The decoder $h$ reconstructs $\mathbf{X}_{k}$ from $(\mathbf{Z}_{k},\mathbf{X}_{\mathcal{B}_{k}})$, yielding $\hat{\mathbf{X}}_{k}=h(\mathbf{Z}_{k},\mathbf{X}_{\mathcal{B}_{k}})$. Under ideal conditions, perfect reconstruction ($\hat{\mathbf{X}}_{k}=\mathbf{X}_{k}$) implies that $h$ effectively approximates the true structural function $f$.
We first elaborate on the architecture and training procedure of PFD-BDCM, and then explain how to employ it for answering causal queries. We start with some notations.
  • Let $Z_{k}^{t}$ be the $k$-th exogenous node value at diffusion step $t$ of the forward implicit diffusion process and let $Z_{k}\coloneqq Z_{k}^{T}$, where $T$ represent the total diffusion steps.
  • Let $\hat{X}_{k}^{t}$ be the $k$-th endogenous node value at diffusion step $t$ of the reverse implicit diffusion process and let $\hat{X}_{k}\coloneqq\hat{X}_{k}^{0}$.
Causal inference, particularly counterfactual reasoning, needs a deterministic mapping from observational data to latent representations (encoding) and subsequent reconstruction (decoding). While standard Denoising Diffusion Probabilistic Models (DDPMs) (Ho et al.,, 2020) employ stochastic reverse processes, Denoising Diffusion Implicit Models (DDIMs) (Song et al.,, 2021) use a non-Markovian formulation that renders the reverse transition deterministic—a property of fundamental importance. Within the PFD-BDCM framework, the encoding and decoding procedures extend the DDIM paradigm by integrating backdoor adjustment sets as supplementary covariates, and the diffusion model for node $k$ is denoted by $\epsilon_{\theta}^{k}(X_{k},\mathbf{X}_{\mathcal{B}_{k}},t)$.
Formally, for each node $k\in[K]$, the latent variable $Z_{k}:=Z_{k}^{T}$ is generated through the forward implicit diffusion process
$Z^{t+1}_{k}$= $\sqrt{\alpha_{t+1}/\alpha_{t}}Z^{t}_{k}+\epsilon_{\theta}^{k}(Z^{t}_{k},\mathbf{X}_{\mathcal{B}_{k}},t)$ (5)
$\times(\sqrt{1-\alpha_{t+1}}-\sqrt{\alpha_{t+1}(1-\alpha_{t})/\alpha_{t}}),$
for $t=0,\cdots,T-1$, initialized with $Z_{k}^{0}:=X_{k}$. This latent representation $Z_{k}$ serves as a proxy for the exogenous noise $U_{k}$. The reconstruction $\hat{X}_{k}:=\hat{X}_{k}^{0}$ is obtained via the reverse implicit diffusion process
$\hat{X}^{t-1}_{k}$= $\sqrt{\alpha_{t-1}/\alpha_{t}}\hat{X}^{t}_{k}-\epsilon_{\theta}^{k}(\hat{X}_{k}^{t},\mathbf{X}_{\mathcal{B}_{k}},t)$ (6)
$\times\left(\sqrt{\alpha_{t-1}(1-\alpha_{t})/\alpha_{t}}-\sqrt{1-\alpha_{t-1}}\right),$
for $t=T,\cdots,1$, initialized with $\hat{X}^{T}_{k}:=Z_{k}$. The encoding and decoding functions for node $k$ are denoted as $\operatorname{Enc}_{k}(X_{k},\mathbf{X}_{\mathcal{B}_{k}})$ (Eq. (5)) and $\operatorname{Dec}_{k}(Z_{k},\mathbf{X}_{\mathcal{B}_{k}})$ (Eq. (6)) respectively, whose pseudocodes are provided in Appendix A.1.
Training PFD-BDCMs.
The comprehensive training methodology (Algorithm 1) incorporates backdoor adjustment sets as covariates while training distinct diffusion models per node. Crucially, generative models for endogenous nodes exhibit mutual independence during training, thereby enabling parallelized optimization. This parallelism is feasible since each diffusion model necessitates only its target node’s values and corresponding backdoor adjustment set values. The final PFD-BDCM architecture integrates these $K$ trained diffusion models $\{\epsilon_{\theta}^{k}\}_{k\in[K]}$.
We now elucidate the methodology for leveraging trained PFD-BDCMs to approximate diverse causal queries. Resolution of observational and interventional queries necessitates sampling from their respective observational and interventional distributions. Counterfactual queries, however, operate at unit granularity by modifying structural equation assignments while preserving the latent exogenous noise variables consistent with empirical observations.
Generating samples for observational/interventional queries. To generate samples approximating the interventional distribution $p(\mathbf{X}|\mathrm{do}(X_{\mathcal{L}}:=\boldsymbol{\gamma}))$ using a trained PFD-BDCM model, we implement the following procedure: i) For intervened nodes $l\in\mathcal{L}$, we set $\hat{X}_{l}:=\gamma_{l}$ deterministically; ii) For root nodes $k$, sample $\hat{X}_{k}$ from empirical training distributions; iii) For non-intervened nodes $k\notin\mathcal{L}$, sample latent vectors $Z_{k}\sim\mathcal{N}(\mathbf{0},I_{d_{k}})$, where $d_{k}=\dim(X_{k})$, and subsequently compute $\hat{X}_{k}:=\mathrm{Dec}_{k}(Z_{k},\hat{\mathbf{X}}_{\mathcal{B}_{k}})$ utilizing inductively generated backdoor variable values $\hat{\mathbf{X}}_{\mathcal{B}_{k}}$. Generated values propagate to child nodes as backdoor inputs. Observational sampling ($p(\mathbf{X})$) corresponds to $\mathcal{L}=\emptyset$, with pseudocode formalized in Algorithm 2.
Algorithm 1 PFD-BDCMs Training
Input: target distribution $\mathcal{Q}$, scale factors $\{\alpha_{t}\}_{t=1}^{T}$, ST-DDAG $\mathcal{G}$ node $k$ is represented by $X_{k}$ and intervened node $l$ with intervened value $\gamma_{l}$}
1:while not converged do
2: Sample $X_{k}^{0}\sim\mathcal{Q}$
3:for $k=1,\cdots,K$ do
4:  $t\sim\mathrm{Unif}[\{1,\cdots,T\}]$, $\varepsilon\sim\mathcal{N}(0,I)$
5:  Update the parameter of the node $k$ ’s diffusion model $\epsilon_{\theta}^{k}$ by minimizing the Eq. (1) depending on the nodes.
6:  if $X_{l}\in\mathbf{X}_{\mathcal{B}_{k}}$ then
7:   $\left\|\epsilon-\epsilon_{\theta}^{k}\left(\sqrt{\alpha_{t}}X_{k}^{0}+\sqrt{1-\alpha_{t}}\epsilon,\mathbf{X}_{\mathcal{B}_{k}}^{0},X_{l},t\right)\right\|_{2}^{2}$
8:  else
9:   $\left\|\epsilon-\epsilon_{\theta}^{k}\left(\sqrt{\alpha_{t}}X_{k}^{0}+\sqrt{1-\alpha_{t}}\epsilon,\mathbf{X}_{\mathcal{B}_{k}}^{0},t\right)\right\|_{2}^{2}$
10:  end if
11:end for
12:end while
Algorithm 2 PFD-BDCMs Observational/Interventional Sampling
Input: Intervened node $l$ with value $\gamma_{l}$ ($\mathcal{L}=\emptyset$ for observational sampling).
1:for $k=1,\cdots,K\{\text{in topological order}\}$  do
2:if  $k=l$  then
3:  $\hat{X}_{k}\leftarrow\gamma_{l}$
4:else if $k$ is a root node then
5:  $\hat{X}_{k}\sim E_{k}$
6:else if $X_{l}\in\mathbf{X}_{\mathcal{B}_{k}}$ then
7:  $\hat{X}_{k}\leftarrow\operatorname{Dec}_{k}(Z_{k},\hat{\mathbf{X}}_{\mathcal{B}_{k}},X_{l})$
8:else
9:  $\hat{X}_{k}\leftarrow\operatorname{Dec}_{k}(Z_{k},\hat{\mathbf{X}}_{\mathcal{B}_{k}})$
10:end if
11:end for
[…]
Counterfactual Queries. To compute counterfactual estimates $\hat{\mathbf{x}}^{\mathrm{CF}}$ within the PFD-BDCM framework, given factual observation $\mathbf{x}^{\mathrm{F}}:=(x_{1}^{\mathrm{F}},\cdots,x_{K}^{\mathrm{F}})$ and intervention set $\mathcal{L}$ with values $\boldsymbol{\gamma}$, we implement the following systematic procedure: i) For intervened nodes $l\in\mathcal{L}$, assign $\hat{x}_{l}^{\mathrm{CF}}:=\gamma_{l}$ deterministically; ii) For non-intervened descendant nodes $k$, using inductively generated backdoor estimates $\hat{\mathbf{x}}_{\mathcal{B}_{k}}^{\mathrm{CF}}$, we compute $\hat{x}_{k}^{\mathrm{CF}}:=\mathrm{Dec}_{k}(\mathsf{Enc}_{k}(x_{k}^{\mathrm{F}},\mathbf{x}_{\mathcal{B}_{k}}^{\mathrm{F}}),\hat{\mathbf{x}}_{\mathcal{B}_{k}}^{\mathrm{CF}})$, where factual noise is implicitly encoded. The complete formalization appears in Appendix A.2.
4 Bounding Counterfactual Error
In this section, we establish the theoretical guarantees for the counterfactual estimation accuracy of the PFD-BDCM framework. The primary contribution is the derivation of an error bound that formally links the reconstruction fidelity of the encoder-decoder architecture to the precision of its counterfactual predictions. Notably, these theoretical guarantees accommodate higher-dimensional settings through additional structural assumptions. Formal proofs are presented in Appendix B.
Consider an endogenous variable $X_{k}$ governed by structural equation $X_{k}:=f_{ij}(X_{\mathcal{B}_{k}},U_{k})$ with backdoor adjustment set $X_{\mathcal{B}_{k}}$ and exogenous noise $U_{k}$. We analyze a single node without loss of generality (by permutation invariance of nodes), henceforth denoting the target variable as $X\in\mathcal{X}\subseteq\mathbb{R}$, its backdoors as $X_{\mathcal{B}}\in\mathcal{X}_{\mathcal{B}}\subseteq\mathbb{R}^{K}$, and exogenous noise as $U$. The encoder-decoder architecture comprises
$g$ $:\mathcal{X}\times\mathcal{X}_{\mathcal{B}}\to\mathcal{Z}\quad\text{(encoding function)},$
$h$ $:\mathcal{Z}\times\mathcal{X}_{\mathcal{B}}\to\mathcal{X}\quad\text{(decoding function)},$
where $\mathcal{Z}$ denotes the latent space. Within PFD-BDCM, $g$ and $h$ correspond to the $\operatorname{Enc}$ and $\operatorname{Dec}$ operators, respectively.
Our theoretical results rely on a set of assumptions regarding the structural equation and the encoder-decoder model. These conditions are essential for ensuring that the latent variable learned by the encoder can uniquely recover the unobserved exogenous noise, which is the cornerstone of accurate counterfactual estimation (Lu et al.,, 2020; Nasr-Esfahany and Kiciman,, 2023; Nasr-Esfahany et al.,, 2023). For a variable $X\in\mathcal{X}\subset\mathbb{R}$ with structural equation $X:=f_{ij}(X_{\mathcal{B}},U)$ where exogenous noise $U\sim\mathcal{N}(0,\psi)$ and $U\perp\!\!\!\perp X_{\mathcal{B}}$, we have the following assumptions:
Assumption 1.
The encoded latent variable is independent of the backdoor variables, $g(X,X_{\mathcal{B}})\perp\!\!\!\perp X_{\mathcal{B}}$.
Assumption 2.
The structural equation $f_{ij}$ is differentiable and strictly increasing with respect to $U$ for all values of the backdoor variables $x_{\mathcal{B}}\in\mathcal{X}_{\mathcal{B}}$.
Assumption 3.
The encoding function $g$ is invertible and differentiable with respect to its first argument $X$ for all $x_{\mathcal{B}}\in\mathcal{X}_{\mathcal{B}}$.
These assumptions, while formal, are well-motivated in the context of causal inference and deep generative models. Assumption 1 ensures that the encoder learns a representation of the exogenous noise that is not contaminated by information from the backdoor variables. This is naturally satisfied in settings like additive noise models with $f_{ij}(X_{\mathcal{B}},U)=f_{ij}^{\prime}(X_{\mathcal{B}})+U$ where $X_{\mathcal{B}}$ and $U$ is independent. If the fitted model $\hat{f}_{ij}\equiv f_{ij}^{\prime}$, then $g(X,X_{\mathcal{B}})=U$. Assumption 2 is satisfied by major identifiable model classes-including additive noise, post-nonlinear, and heteroscedastic formulations (Strobl and Lasko,, 2023) while concurrently resolving symmetric noise ambiguities characteristic of observational data. This assumption further aligns with contemporary identifiability frameworks (Nasr-Esfahany and Kiciman,, 2023) and intrinsically precludes non-identifiable structural equations. The encoder invertibility condition (Assumption 3) is intrinsically satisfied by the bijective properties of deterministic diffusion architectures (Song et al.,, 2021), guaranteeing uniqueness in latent representations while preserving compatibility with standard implementations.
Under these assumptions, we can prove that the encoder successfully isolates the exogenous noise up to an invertible transformation.
Theorem 1.
Under Assumptions 1, 2, and 3, the encoded latent variable $Z=g(X,X_{\mathcal{B}})$ is an invertible transformation of the true exogenous noise $U$. That is, there exists an invertible function $\tilde{q}$ such that $Z=\tilde{q}(U)$.
This theorem provides the foundation for assessing the accuracy of counterfactuals given by the PFD-BDCM. It implies that the abduction step, $\operatorname{Enc}(X,X_{\mathcal{B}})$, correctly captures the essence of the unobserved confounder $u$ that generated the factual observation. We now explore the direct consequences of this result.
Perfect Estimation.
In an oracle scenario where the model achieves perfect reconstruction which means $h(g(X,X_{\mathcal{B}}),X_{\mathcal{B}})=X$ holding almost surely (a.s.), the counterfactual estimate will be “perfect”. This precise case implies the satisfaction of Theorem 1, yielding the relationship $h(\tilde{q}(U),X_{\mathcal{B}})=f_{ij}(X_{\mathcal{B}},U)$. Consequently, when making a counterfactual prediction for a new intervention $X_{\mathcal{B}}:=\gamma$, the model computes $h(\tilde{q}(u),\gamma)$, which a.s. equates to the true counterfactual $f_{ij}(\gamma,u)$.
Corollary 1.
Assume Assumptions 1, 2, and 3 hold and the encoder-decoder model pair $(g,h)$ achieves perfect reconstruction, i.e., $h(g(X,X_{\mathcal{B}}),X_{\mathcal{B}})=X$ a.s.. For a factual observation pair $(x,x_{\mathcal{B}})$ generated by $x=f_{ij}(x_{\mathcal{B}},u)$ and a counterfactual intervention $\mathrm{do}(X_{\mathcal{B}}:=\gamma)$, the estimated counterfactual $\hat{x}^{\mathrm{CF}}:=h(g(x,x_{\mathcal{B}}),\gamma)$ is a.s. identical to the true counterfactual $x^{\mathrm{CF}}:=f_{ij}(\gamma,u)$.
More practically, models are not perfect. Corollary 2 allows us to bound the counterfactual error by the model’s reconstruction error. This is a powerful result, as it connects a measurable property of the model (how well it auto-encodes data) to its performance on a causal task.
Corollary 2.
Under Assumptions 1, 2, and 3, if the model’s reconstruction error is bounded by $\tau$ under a metric $d$, such that $d(h(g(X,X_{\mathcal{B}}),X_{\mathcal{B}}),X)\leq\tau$, then the counterfactual estimation error is also bounded by $\tau$. For a factual observation $(x,x_{\mathcal{B}})$ and intervention $\mathrm{do}(X_{\mathcal{B}}:=\gamma)$, we have $d(\hat{x}^{\mathrm{CF}},x^{\mathrm{CF}})\leq\tau$.
This corollary formally establishes that minimizing the reconstruction loss during training directly optimizes the model for better counterfactual prediction.
The framework extends to the more general multivariate setting where $X\in\mathbb{R}^{d}$. This requires a stronger assumption on the encoder’s Jacobian to ensure that information is not lost in the higher-dimensional space.
Theorem 2.
Consider $X\in\mathcal{X}\subset\mathbb{R}^{d}$ and $U\in\mathbb{R}^{m}$ with $m\geq d$. Assume the following conditions:
  1. 1.
    The structural equation $X=f_{ij}(X_{\mathcal{B}},U)$ is bijective in $U$ for any fixed $X_{\mathcal{B}}$.
  2. 2.
    The encoder’s Jacobian, $\partial g/\partial X$, has full rank $d$ almost everywhere.
  3. 3.
    The decoder $h(Z,X_{\mathcal{B}})$ is $L_{h}$-Lipschitz with respect to its first argument $Z$.
  4. 4.
    The reconstruction error is bounded such that $\|h(g(X,X_{\mathcal{B}}),X_{\mathcal{B}})-X\|\leq\tau$.
Then for a counterfactual intervention $\mathrm{do}(X_{\mathcal{B}}:=\gamma)$, the estimation error is bounded by:
$\|\hat{x}^{\mathrm{CF}}-x^{\mathrm{CF}}\|\leq L_{h}\cdot\kappa_{g}\cdot\tau,$
where $\kappa_{g}=\sup\|(\partial g/\partial X)^{-1}\|_{2}$ is the bound for the inverse of the encoder’s Jacobian.
5 Experimental Evaluation
We empirically evaluated the efficacy of PFD-BDCM in addressing causal queries across synthetic and real-world datasets. All experiments were implemented in $\mathrm{Python}$ (version 3.9.0) on a Windows server equipped with an Intel i9-13900K processor, NVIDIA RTX 4090 24G GPU, and 128 GB RAM. To demonstrate that PFD-BDCM faithfully samples from the target interventional distribution, we designed scenarios where causal sufficiency was deliberately violated within partially functional structural models.
5.1 Simulation Study
Figure 2 depicted in Appendix C.1 presented the instantiated PFST-DSCM, where causal sufficiency was compromised. Consider $\{X_{28},X_{29}\}$ as unobserved explanatory variables and $X_{30}$ as an unobserved explained variable, we assumed exhibit pronounced spatial heterogeneity coupled with temporal dependence between $\{X_{28},X_{29}\}$ and $X_{30}$. Let $X_{1},X_{2},X_{3}$ represent endogenous cause variables; $X_{31},X_{32},X_{33}$ denote outcome variables; and $\boldsymbol{X}_{\mathcal{B}}=\{X_{22},\cdots,X_{27}\}$ constitute backdoor adjustment sets. For visual clarity, exogenous noise terms $\boldsymbol{U}$ were omitted.
The partial functional dynamic structural equations were defined as: $X_{ijk}=f_{ij}(X_{\mathcal{B}k},U_{k})$ (Eq. (23) in Appendix C.1). The structural equations governing the Partially Functional Dynamic Diffusion-based Causal Model(PFD-DCM) and PFD-BDCM were instantiated with additive noise models (ANM) (Peters et al.,, 2013) furnishing elementary baselines.
Our objective was to accurately sample from the post-interventional distribution $q(X_{k}|\operatorname{do}(X_{l}=\gamma_{l}))$, where $k\in\{31,32,33\}$ indexes outcomes and $l\in\{1,2,3\}$ indexes causes. During intervention, $X_{l}$ is fixed to $\gamma_{l}$, while root variables $X_{h}$ ($h=1,\cdots,21$) were sampled from their empirical marginals $E_{h}$. For outcome $X_{k}$, PFD-DCM employed $\operatorname{Dec}_{k}(Z_{k},\hat{X}_{h})$, whereas PFD-BDCM utilized $\operatorname{Dec}_{k}(Z_{k},\hat{X}_{\mathcal{B}_{k}},\hat{X}_{h})$, thereby leveraging backdoor adjustments. Comprehensive simulation setting was detailed in the Appendix C.1.
Table 1 summarizes aggregated performance metrics-observational (Obs.), interventional (Int.), and counterfactual (CF.)-averaged over nine independent random initializations. Comprehensive diagnostics (boxplots and kernel density estimates) were provided in Appendix C.1. PFD-DCM and PFD-BDCM achieved compelling statistical fidelity across all query types, evidenced by MMD and CRPS for observational/interventional query and low MSE for counterfactual query. Crucially, PFD-BDCM consistently outperformed PFD-DCM, attributable to its principled utilization of backdoor adjustments that mitigated information loss from critical latent nodes.
$J=6$ PFD-DCM PFD-BDCM
$n=30$ $3.925\pm 3.792$ $3.532\pm 3.920$
[…]
(Obs.) $n=200$ $0.737\pm 0.672$ $0.533\pm 0.486$
$n=500$ $0.263\pm 0.250$ $0.201\pm 0.136$
$n=30$ $0.559\pm 0.070$ $0.552\pm 0.070$
[…]
(Obs.) $n=200$ $0.558\pm 0.025$ $0.556\pm 0.024$
$n=500$ $0.565\pm 0.015$ $0.563\pm 0.015$
$n=30$ $2.705$ $2.665$
[…]
(Obs.) $n=200$ $15.309$ $15.671$
$n=500$ $44.661$ $44.918$
$n=30$ $4.021\pm 3.749$ $4.298\pm 3.946$
[…]
(Int.) $n=200$ $0.558\pm 0.457$ $0.579\pm 0.530$
$n=500$ $0.245\pm 0.207$ $0.217\pm 0.203$
$n=30$ $0.557\pm 0.072$ $0.557\pm 0.067$
[…]
(Int.) $n=200$ $0.562\pm 0.025$ $0.561\pm 0.027$
$n=500$ $0.565\pm 0.017$ $0.562\pm 0.016$
$n=30$ $2.070$ $2.191$
[…]
(Int.) $n=200$ $13.565$ $13.584$
$n=500$ $38.233$ $38.299$
$n=30$ $0.584\pm 0.212$ $0.815\pm 0.252$
[…]
(CF.) $n=200$ $0.074\pm 0.020$ $0.598\pm 0.095$
$n=500$ $0.024\pm 0.006$ $0.579\pm 0.056$
$n=30$ $1.537$ $1.535$
[…]
(CF.) $n=200$ $8.621$ $8.616$
$n=500$ $22.163$ $22.167$
[…]
5.2 Empirical Application
This investigation employed the PFD-BDCM framework to examine spatio-temporal dynamic structural causal relationships among air pollutant indicators and their determinants. Our analysis encompassed 30 provincial-level administrative divisions across Chinese mainland during the period January 2015 to December 2020. The study integrates China’s provincial CO2 emission inventories from the China Emission Accounts and Datasets (CEADs) (Shan et al.,, 2018, 2020; Guan et al.,, 2021; Xu et al.,, 2024) and emissions data for nine atmospheric pollutants from the Multi-scale Emission Inventory of China (MEIC) (Li et al.,, 2018; Zheng et al.,, 2018; Geng et al.,, 2024; Li et al.,, 2019) as response variables for air pollutant emissions.
Building upon prior research (Ghosh,, 2010; Jayanthakumaran et al.,, 2012; Ozcan,, 2013; Zhu et al.,, 2021) and incorporating domain-specific characteristics of regional emissions, we systematically collected foundational determinants across ten conceptual dimensions. The comprehensive dataset comprised 118 indicator variables, through collinearity diagnostics and random forest-based feature selection, we retained 49 statistically robust indicators for subsequent modeling (detailed indicators show in Appendix C.2 Table 6).
Complete experimental specifications and supplementary materials were documented in Appendix C.2, with observational query results presented in Table 2 .
Variable PFD-DCM PFD-BDCM
SO2 $0.495\pm 0.419$ $0.512\pm 0.476$
NOx $0.511\pm 0.449$ $0.473\pm 0.424$
CO $0.564\pm 0.531$ $0.524\pm 0.450$
VOC $0.489\pm 0.464$ $0.480\pm 0.403$
NH3 $0.519\pm 0.538$ $0.494\pm 0.446$
PM10 $0.487\pm 0.426$ $0.485\pm 0.457$
PM2.5 $0.400\pm 0.356$ $0.435\pm 0.400$
BC $0.417\pm 0.406$ $0.350\pm 0.339$
OC $0.487\pm 0.462$ $0.398\pm 0.413$
CO2 $0.336\pm 0.257$ $0.357\pm 0.304$
[…]
6 Concluding Remarks
This research introduces the Partial Functional Dynamic Backdoor Diffusion-based Causal Model (PFD-BDCM), a novel methodological framework crafted for robust causal inference amidst spatial heterogeneity, temporal dependencies, and unmeasured confounding. Our principal contributions are threefold.
Model Innovation: PFD-BDCM synergistically integrates functional basis expansions with diffusion-based causal modeling, facilitating simultaneous resolution of: i) Multi-resolution variables through partial functional representations; ii) Spatio-temporal dynamics via regionally parameterized structural equations; iii) Unmeasured confounder bias utilizing backdoor adjustment sets.
Theoretical Foundation: We establish pioneering error bounds formally connecting counterfactual estimation accuracy to encoder-decoder reconstruction fidelity under: i) Monotonic structural functional constraints; ii) Invertible encoding operators; iii)Multivariate generalizations with supplementary structural assumptions.
Empirical Validation: Comprehensive experiments on synthetic and real-world data demonstrate that PFD-BDCM significantly outperforms existing methods in answering observational, interventional, and counterfactual queries.
While PFD-BDCM advances causal inference in complex settings, future work should address: i) Scalability enhancements for ultra-high-dimensional functional data via tensor decompositionsp; ii)Automated backdoor set identification through causal discovery algorithms; iii)Temporal graph neural network integration for non-stationary processes; iv) Real-time deployment in environmental policy decision support systems.
The proposed framework opens new avenues for causal inference in environmental science, epidemiology, and econometrics where functional data and unmeasured confounders are prevalent.